OpenStudy (anonymous):

Having trouble with another if someone has the time to help me. I would really appreciate it. I need to figure out how to solve this equation: 3/4 = 24y and if you could show me how you come to the solution, which should be 1/32. Thanks in advance for any help, Denise

OpenStudy (anonymous):

you're solving for y so you want to get y by itself. right now it is multiplied by 24 so divide both sides by 24. Dividing by 24 is the same as multiplying by 1/24. Our equation then becomes y=(3/4)*(1/24). Multiply to get y=3/(24*4). The denominator has a 3 factor which cancels with the numerator to get the result.

OpenStudy (anonymous):

Thanks Igrim, it was the rule about dividing I totally spaced. Thank you so much, excellent help, Denise!!

OpenStudy (anonymous):

why isn't it 24/1 though? I mean (3/4) x (24/1)?

OpenStudy (anonymous):

Ok, I see now why, in your above note. sorry I didn't get it the first time;)

OpenStudy (anonymous):

you're dividing by 24. 3/4*24/1 would be multiplying both sides by 24. This might help: we started with 3/4 = 24y. We want to move the 24 so if we multiply by 1/24, our equation becomes (1/24)* 3/4 = (1/24)*24y. But (1/24)*24y=24/24y=y. Okay. I'm glad you got it.

OpenStudy (anonymous):

Ok, thank you for your time!! Denise:)